OpenStudy (anonymous):

The ratio of the measures of two complementary angles is 1:5. What are the measures of the angles?

OpenStudy (anonymous):

A. 30° and 60° B. 15° and 75° C. 30° and 150° D. 5° and 25

OpenStudy (anonymous):

complementary angles always add up to 180 degrees

OpenStudy (anonymous):

Thanks

OpenStudy (raden):

NO, add up to 90 degrees. supplementary is add up to 180 degrees!

OpenStudy (anonymous):

So the answer would be C?

OpenStudy (anonymous):

lol

OpenStudy (raden):

1+5 =6 1st angles = 1/6 * 90 = 15 2nd angle = 5/6 * 90 = 75

OpenStudy (anonymous):

Oops, @RadEn is right about complementary adding up to 90